OpenStudy (anonymous):

Slove for X: x^4 = 1/625

OpenStudy (anonymous):

\[\sqrt[n]{x^n}=x\]

OpenStudy (anonymous):

take the fourth route of the fraction so (1/625) ^ (1/4) = 1 / (625^(1/4)) = 1/5

jimthompson5910 (jim_thompson5910):

Here's how I would do it x^4 = 1/625 625x^4 = 1 625x^4 - 1 = 0 (25x^2)^2 - 1^2 = 0 (25x^2 - 1)(25x^2 + 1) = 0 ((5x)^2 - 1^2)(25x^2 + 1) = 0 (5x - 1)(5x + 1)(25x^2 + 1) = 0 Do you see how to finish up?

OpenStudy (anonymous):

\[\sqrt[4]{x ^{4}}=\sqrt[4]{1/625}\] Like this easy ?.. :)